OpenStudy (anonymous):

The consumer demand equation for tissues is given by: q = (106 − p)^2, where p is the price per case of tissues and q is the demand in weekly sales. (a) Determine the price elasticity of demand E when the price is set at $32. (Round your answer to three decimal places.) E = ____? The demand is going down by what percent per 1% increase in price at that price level? (b) At what price should tissues be sold in order to maximize the revenue? (Round your answer to the nearest cent.) $_____? (c) Approximately how many cases of tissues would be demanded at that price? ____ cases per week?

OpenStudy (anonymous):

got to love economics :)

OpenStudy (anonymous):

hmmmm the elasticity would be the change in quantity demanded divided by the change in price

OpenStudy (anonymous):

but i see no original price or quantity

OpenStudy (anonymous):

quantity = 5476 at $32

OpenStudy (anonymous):

5476/32 = 171.188

OpenStudy (anonymous):

so a. e= 171.188 i believe

OpenStudy (anonymous):

1% of $32 is $.32

OpenStudy (anonymous):

5476/32.32 = 177.785

OpenStudy (anonymous):

5476/32.64 = 167.785

OpenStudy (anonymous):

5476/32.32 = 169.431 oops